OpenStudy (anonymous):

I WILL FAN AND MEDAL!!! one and one fourth A. n over 5=1/4 B. 1/2n=4/5 C. 3/4n=1/8 D. n/16=5/8

OpenStudy (anonymous):

A. n over 5=1/4 B. 1/2n=4/5 C. 3/4n=1/8 D. n/16=5/8

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (anonymous):

what is the question?

OpenStudy (anonymous):

what is the eqation to 1 1/4

OpenStudy (setsuna-yuregeshi):

So, what equation equals1 and 1/4?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

that's why I'm here

OpenStudy (setsuna-yuregeshi):

No, that's not what I meant. I'm asking: Is the question : Which equation equals 1 and 1/4? What is the question asking for?

OpenStudy (anonymous):

yes

OpenStudy (setsuna-yuregeshi):

So, what do you think you need to do?

OpenStudy (anonymous):

IDK!!

OpenStudy (setsuna-yuregeshi):

Do you know how to solve for x? Try that with each one....

OpenStudy (setsuna-yuregeshi):

But start with A,

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i can't

OpenStudy (setsuna-yuregeshi):

You don't know how to solve for x? Or?

OpenStudy (anonymous):

no

OpenStudy (setsuna-yuregeshi):

What math course are you in?

OpenStudy (anonymous):

7th grade

OpenStudy (anonymous):

@Teddyiswatshecallsme

OpenStudy (setsuna-yuregeshi):

So pre algebra?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i hate math